# Fully-Polarized Topological Isostatic Metamaterials in Three Dimensions

### v1

@@ Abstract
In this work, we demonstrate, both theoretically and experimentally, the fully-polarized topological mechanical phase in 3D. Using the three-dimensional example known as the generalized pyrochlore lattice, we illustrate this topological mechanical phase and the resulting distinctive boundary elasticity. Our analytic design principle is based on the mechanical transfer matrix Zhou2018PRL; Zhou et al. (2019) that polarizes all floppy modes to concentrate on a single open boundary of the lattice, whereas the opposite surface is clear of floppy modes. Consequently, the lattice is topologically fully-polarized and exhibits highly contrasting boundary mechanics.
@@
It is worth emphasizing that our 3D topological lattice is fundamentally distinct from its 2D counterparts, because 2D mechanical structures are prone to deform, distort, and lose mechanical stability due to external pressure, thermal expansion, and bending Charara et al. (2022); Zunker and Gonella (2021). Our work discovers the fully-polarized topological mechanical phase in 3D. This new phase opens up avenues towards novel applications of topological mechanical metamaterials, including asymmetric wave propagation, directional polar elasticity in isostatic media, and topological fracturing protection pyr .

### v2

@@ Abstract
In this work, we demonstrate, both theoretically and experimentally, the fully-polarized topological mechanical phase in 3D. Using the three-dimensional example known as the generalized pyrochlore lattice, we illustrate this topological mechanical phase and the resulting distinctive boundary elasticity. Our analytic design principle is based on the mechanical transfer matrix Zhou et al. (2018, 2019) that polarizes all floppy modes to concentrate on a single open boundary of the lattice, whereas the opposite surface is clear of floppy modes. Consequently, the lattice is topologically fully-polarized and exhibits highly contrasting boundary mechanics.
@@
It is worth emphasizing that our 3D topological lattice is fundamentally distinct from its 2D counterparts, because 2D mechanical structures are prone to deform, distort, and lose mechanical stability due to external pressure, thermal expansion, and bending Charara et al. (2022); Zunker and Gonella (2021a). Our work discovers the fully-polarized topological mechanical phase in 3D. This new phase opens up avenues towards novel applications of topological mechanical metamaterials, including asymmetric wave propagation, directional polar elasticity in isostatic media, and topological fracturing protection pyr .